The third term of an AP is 45 and the 23 term term is 315. find the 59th term.
Answer by greenestamps(13200)


You could use the standard formulas for working with arithmetic progressions to solve the problem. However, I believe you will understand what you are doing better (and get to the answer a lot faster!) if you work in a less formal manner.

The 23rd term is 20 terms after the 3rd term; and it is 270 greater than the 3rd term. So the common difference is 270/20 = 13.5.

The 59th term is 36 terms after the 23rd term; so the difference between the 23rd term and the 59th term is 36*13.5 = 486.

So the 59th term is 315+486 = 801.